# E-3DGS: Gaussian Splatting with Exposure and Motion Events

### v1

@@ III-D EME-3D: Exposure and Motion Events Dataset for 3D Reconstruction
We utilize an AT-DVS (Adjustable Transmittance DVS) to capture real-world sequences. Fig. 2 illustrates the experimental setup used for data acquisition. The optical motorized rotating stage in our setup offers a high precision of $0.005$°. Before the experiment, the object is securely affixed to the stage using Blu-Tack to prevent any relative movement during rotation. Each object undergoes two recording phases, with the rotating stage reset to the zero position before each phase. In the first phase, the stage performs a full $360$° rotation to capture motion events. During the second phase, the stage rotates in $1.8$° increments, controlled by a motion controller, to capture exposure events at specific positions.
TABLE II: Qualitative Comparisons of neural reconstruction.
@@
IV Experiments
IV-A Implementation Details
Our implementation is based on the official 3D-GS codebase [10]. For training in the fast reconstruction mode, which uses only motion event supervision, we run for $10,000$ iterations with a $500$-iteration warmup and an initial learning rate of $1.6e^{-5}$. For high-quality reconstruction mode and balanced hybrid mode, we run for $30,000$ iterations with a $500$-iteration warmup and an initial learning rate of $1.6e^{-4}$. All experiments were performed on a single NVIDIA RTX 3090-Ti GPU.
IV-B Baselines
We benchmark our approach against a NeRF-based method, EventNeRF [14], and a naive baseline, E2VID [20] + NeRF [9], which cascades the event-to-video pipeline E2VID into NeRF. Additionally, we compare a method that cascades E2VID with a vanilla 3D Gaussian Splatting approach [10]. We observed significant background noise in the images reconstructed by the learning-based E2VID, which adversely affects the accuracy of E2VID + 3DGS. To address this, we have introduced a stronger baseline by introducing the Segment Anything Model (SAM) [32] to segment and remove background noise from the grayscale images reconstructed by E2VID. The segmentation quality was manually double-checked and refined, forming our strongest baseline: E2VID + SAM + 3DGS. For real-world scenes, we retrained EventNeRF and all baselines on the EME-3D dataset, rendering RGB and depth images for comparison to ensure fairness and reliability.
IV-C Quantitative Comparison
As shown in Tab. II, we compare the proposed E-3DGS method with other baselines (see Sec. IV-B) on the real-world EME-3D dataset. In fast reconstruction mode, E-3DGS outperforms all baselines across all sequences in terms of reconstruction quality. On average, E-3DGS (Fast Reconstruction Mode), using only motion events, achieves a PSNR improvement of $2.19dB$ ($25.22$ vs. $23.03$) over the strongest baseline E2VID + SAM + 3DGS, which utilizes additional learning-based events-to-video [20] and vision foundation model [32] techniques. Moreover, compared to the classic event-only method EventNeRF [14], E-3DGS (Fast Reconstruction Mode) shows a PSNR improvement of $5.68dB$ ($25.22$ vs. $19.54$). In high-quality reconstruction mode, E-3DGS fully leverages the high-resolution spatial cues from exposure events to optimize the Gaussian ellipsoids and distributions, resulting in a substantial further improvement in reconstruction quality. On average, E-3DGS achieves a PSNR increase of $10.89dB$ ($33.92$ vs. $23.03$), an SSIM improvement of $0.08$ ($0.97$ vs. $0.89$), and an LPIPS reduction of $0.03$ ($0.09$ vs. $0.12$) compared to the strongest E2VID + SAM + 3DGS baseline. Notably, the training process of E-3DGS does not rely on the complex computations required by the learning-based E2VID or SAM, and E-3DGS also significantly outperforms the EventNeRF family in rendering speed ($79.37$ FPS vs. $0.03$ FPS).
@@ IV-C Quantitative Comparison
Figure 3: Qualitative comparison of appearance reconstruction on the real-world EME-3D dataset. The E2VID + NeRF method successfully reconstructs the overall scene but lacks fine local details, such as the tortoise’s shell texture. EventNeRF exhibits noticeable artifacts in both the car and the tortoise. In contrast, our proposed E-3DGS, in both fast and high-quality reconstruction modes, preserves sharper and more consistent structures, while maintaining cleaner backgrounds. Moreover, compared to E2VID + 3DGS, E-3DGS leveraging exposure events demonstrates superior detail and brightness reconstruction.
@@
Figure 4: Qualitative comparison of geometry reconstruction on the real-world EME-3D dataset. Both E2VID+NeRF and EventNeRF, which are based on NeRF, struggle to separate the foreground from the background and are affected by noticeable noise. In contrast, the physics-based 3DGS method handles geometry reconstruction more effectively. Compared to E2VID+3DGS, our E-3DGS excels in capturing high-frequency spatial details, such as the crab’s body and the texture of the base.
IV-D Qualitative Comparison
@@
We qualitatively compare the appearance and geometric reconstruction of five sequences from the real-world EME-3D dataset by visualizing the rendered results and depth maps. Fig. 3 shows that our E-3DGS, in both fast and high-quality reconstruction modes, preserves sharper, more consistent structures and cleaner backgrounds compared to EventNeRF [14]. When using motion events, E-3DGS avoids artifacts seen in EventNeRF’s rendering, such as with the car ($1^{st}$ row) and the tortoise ($4^{th}$ row). Our results also exhibit higher contrast and sharper spatial details, particularly in regions with highlights and reflections, like the camera ($3^{rd}$ row) and the tortoise ($4^{th}$ row). Compared to E2VID [20] + NeRF, E-3DGS produces clearer local details, such as the tortoise’s shell texture, which is completely lost in the $4^{th}$ row of E2VID + NeRF. Furthermore, compared to E2VID + 3DGS, E-3DGS integrates exposure event information and reconstructs overall colors and brightness more accurately. For instance, in the $3^{rd}$ row, E2VID + 3DGS shows white artifacts on the black camera body, while E-3DGS renders a clean, pure black camera body. As shown in Fig. 4, E-3DGS achieves cleaner backgrounds, fewer artifacts, and richer spatial details in geometric reconstruction compared to both EventNeRF and EventNeRF integrated with E2VID. EventNeRF’s depth maps struggle to distinguish foreground from background and suffer from noticeable noise, such as with the car ($1^{st}$ row), motorcycle ($2^{nd}$ row), tortoise ($4^{th}$ row), and crab ($5^{th}$ row). While the depth map quality of E2VID + 3DGS is comparable to E-3DGS, E-3DGS excels in high-frequency spatial details. For example, in the crab sample, E-3DGS successfully reconstructs the gap between the crab’s body and the base, while E2VID + 3DGS confuses the subject with the background.
IV-E Ablations
@@
Figure 5: Qualitative comparison of different initialization methods.
TABLE III: Impact of Initialization Methods on Fast and HQ Modes.
@@
Point Cloud Initialization. In Tab. III, we explore the impact of different point cloud initialization methods on rendering performance in event-based 3DGS. Compared to random initialization, using a sparse point cloud from SfM [12] significantly improves the rendering accuracy when using motion events, with PSNR elevating to $25.22dB$ from $15.72dB$. However, in high-quality reconstruction mode, the PSNR is not sensitive to whether the initialization is from SfM or random point clouds ($33.82dB$ vs. $33.92dB$). As shown in Fig. 5, we visualize the effects of different initialization methods on event-based 3DGS rendering. When using random initialization, motion event-based 3DGS produces noticeable artifacts and limited details (top-left vs. bottom-left). In contrast, after integrating the spatially rich exposure events, E-3DGS becomes more robust to the initialization method, with no significant differences in rendering results (top-right vs. bottom-right). To further illustrate the advantages of incorporating exposure events in 3D reconstruction, we compare NeRF-based and event-to-video-based point cloud initialization methods in Tab. IV. Initializing the point cloud using EventNeRF-rendered results, compared to random initialization, provides a modest improvement in PSNR ($17.13dB$ vs. $15.72dB$), though it requires additional training for EventNeRF with limited gains. Using E2VID [20] to convert events into images followed by SfM for point cloud initialization further enhances accuracy ($23.47dB$ vs. $15.72dB$), but this process also involves additional learning-based computation. By fully leveraging the rich spatial cues embedded in exposure events, the proposed exposure event temporal-spatial mapping combined with SfM initialization yields the greatest improvement in reconstruction quality ($25.22dB$ vs. $15.72dB$). This certifies the critical role of exposure events in enhancing the quality of explicit 3D reconstruction from events.
TABLE IV: Ablation Study on Different Initialization Method.
@@
Figure 6: Effects of integrating motion events with exposure events on 3D reconstruction for the real-world car sequence.
@@
TABLE V: Quantitative results of novel view reconstruction.
@@
Novel View Synthesis. In this task, we trained the model using half of the odd-numbered frames (Given) and tested its ability to generate novel views using the remaining even-numbered frames (Novel). As shown in Tab. V, E-3DGS has excellent novel view synthesis performance in both fast reconstruction mode and high-quality reconstruction mode. On average, E-3DGS using motion events exhibits only a $0.05dB$ drop in PSNR for novel view synthesis compared to the given frames ($23.04dB$ vs. $23.09dB$). Similarly, E-3DGS using exposure events shows a $0.06dB$ PSNR drop for novel views compared to the given frames ($24.83dB$ vs. $24.89dB$). This demonstrates E-3DGS’s strong capabilities in both 3D reconstruction and novel view synthesis tasks.
V Conclusion
@@
In this work, we have presented E-3DGS, the first method to integrate 3DGS with exposure events for explicit scene reconstruction using a single pure high-resolution event sensor. By partitioning events into motion and exposure categories, E-3DGS leverages motion events for efficient 3D Gaussian Splatting (3DGS) reconstruction and utilizes exposure events to enhance reconstruction quality through high-resolution temporal-to-intensity mapping. Our framework is versatile, capable of operating on motion events alone or adopting a hybrid mode that balances reconstruction quality and speed by combining initial exposure events with high-speed motion events. Additionally, we introduced EME-3D, a real-world dataset with exposure and motion events, camera calibration parameters, and sparse point clouds to support further research in this domain. E-3DGS outperforms event-based NeRF in both reconstruction speed and quality while being more cost-effective than methods that require both event and RGB data. By fully utilizing both motion and exposure events, E-3DGS sets a new benchmark for event-based 3D reconstruction, demonstrating robust performance in challenging conditions with reduced hardware demands. In the future, we intend to explore the application and enhancement of E-3DGS based on exposure events for dense robotic perception, aiming to achieve higher-quality large-scale dense 3D reconstruction using purely event-based data.

### v2

@@ 4.1 Evaluation Dataset
4.2 Implementation Details
Our implementation is based on the official 3D-GS codebase [9]. For training in the fast reconstruction mode, which relies solely on motion event supervision, we limit the training schedule to $10,000$ iterations to mitigate the risk of overfitting caused by the sparse spatial detail of motion events. This training includes a $500$-iteration warmup phase with an initial learning rate of $1.6\times 10^{-5}$. For the high-quality reconstruction mode (which utilizes only exposure events for supervision) and the balanced hybrid mode (where supervision transitions from exposure events during slow motion to motion events during fast motion), we run for $30,000$ iterations with a $500$-iteration warm-up and an initial learning rate of $1.6\times 10^{-4}$. In the EventNeRF dataset [12], the ground truth for quantitative evaluation is derived by directly converting the RGB images provided in the dataset to grayscale. For our real-world dataset, the grayscale images mapped from exposure events are used as the ground truth for evaluation. All experiments were performed on a single NVIDIA RTX 3090-Ti GPU.
4.3 Baselines
We benchmark our approach against a NeRF-based method, EventNeRF [12], and a naive baseline, E2VID [22] + NeRF [8], which cascades the event-to-video pipeline E2VID into NeRF. Additionally, we compare a method that cascades E2VID with a vanilla 3D Gaussian Splatting approach [9]. We observed significant background noise in the images reconstructed by the learning-based E2VID, which adversely affects the accuracy of E2VID + 3DGS. To address this, we have introduced a stronger baseline by introducing the Segment Anything Model (SAM) [33] to segment and remove background noise from the grayscale images reconstructed by E2VID. The segmentation quality was manually double-checked and refined, forming our strongest baseline: E2VID + SAM + 3DGS. For the EventNeRF [12] dataset, we convert the color results provided by EventNeRF [12] to grayscale directly for comparison. For our collected real-world scenes, we retrained EventNeRF and all baselines on the EME-3D dataset, rendering RGB and depth images for comparison to ensure fairness and reliability.
4.4 Quantitative Comparison
To evaluate the feasibility of utilizing exposure events for 3D reconstruction, as detailed in Section 4.1, we first simulate exposure events from RGB images and subsequently generate grayscale images using the Temporal-to-Intensity Mapping method. As shown in Tab. 2, leveraging exposure events under the 3DGS framework increases PSNR from $17.29$ (E2VID + 3DGS) to $35.73$. This demonstrates the advantages of hardware-augmented event cameras in generating intensity images over learning-based methods. Across the first six sequences, 3D reconstruction driven by exposure events consistently surpasses the other three approaches in PSNR, SSIM, and LPIPS, demonstrating its robustness and effectiveness. In the mic sequence, EventNeRF [12] slightly outperforms our method due to segmentation errors during exposure event simulation. EfficientSAM [33] faced challenges as the mic sequence’s small foreground and its similarity to the background increased segmentation errors, degrading grayscale reconstruction. Nonetheless, our results highlight the potential of exposure events to enhance event cameras by leveraging temporal intensity dynamics for robust 3D reconstruction.
As shown in Tab. 3, we further compare the proposed E-3DGS method with other baselines (see Sec. 4.3) on the real-world EME-3D dataset. In fast reconstruction mode, E-3DGS outperforms all baselines across all sequences in terms of reconstruction quality. On average, E-3DGS (Fast Reconstruction Mode), using only motion events, achieves a PSNR improvement of $2.19dB$ ($25.22$ vs. $23.03$) over the strongest baseline E2VID + SAM + 3DGS, which utilizes additional learning-based events-to-video [22] and vision foundation model [33] techniques. Moreover, compared to the classic event-only method EventNeRF [12], E-3DGS (Fast Reconstruction Mode) shows a PSNR improvement of $5.68dB$ ($25.22$ vs. $19.54$). In high-quality reconstruction mode, E-3DGS fully leverages the high-resolution spatial cues from exposure events to optimize the Gaussian ellipsoids and distributions, resulting in a substantial further improvement in reconstruction quality. On average, E-3DGS achieves a PSNR increase of $10.89dB$ ($33.92$ vs. $23.03$), an SSIM improvement of $0.08$ ($0.97$ vs. $0.89$), and an LPIPS reduction of $0.03$ ($0.09$ vs. $0.12$) compared to the strongest E2VID + SAM + 3DGS baseline. Notably, the training process of E-3DGS does not rely on the complex computations required by the learning-based E2VID or SAM. Additionally, E-3DGS demonstrates a significant advantage in rendering speed achieving $79.37$ FPS, whereas EventNeRF renders at only $0.03$ FPS (our evaluation) or $0.045$ FPS (as reported in [20] as a baseline comparison).
Table 2: Quantitative Comparisons on the Synthetic Dataset.
Synthetic EventNeRF Dataset Metric Method chair drums ficus hotdog lego materials mic Average PSNR (dB) E2VID [22, 36]+NeRF [8] 18.00 14.01 17.59 23.25 15.75 16.79 15.91 17.33 EventNeRF [12] 29.06 26.91 31.80 29.97 23.17 21.30 31.40 27.66 E2VID [22, 36]+3DGS [9] 17.85 14.47 17.27 23.17 15.68 16.77 15.82 17.29 Ours (Exposure Events) 40.87 34.02 40.00 40.08 27.83 35.96 31.36 35.73 SSIM $\uparrow$ E2VID [22, 36]+NeRF [8] 0.92 0.83 0.91 0.94 0.81 0.90 0.92 0.89 EventNeRF [12] 0.94 0.92 0.94 0.95 0.90 0.94 0.97 0.94 E2VID [22, 36]+3DGS [9] 0.92 0.87 0.91 0.94 0.84 0.91 0.93 0.90 Ours (Exposure Events) 0.99 0.98 0.99 0.99 0.95 0.99 0.97 0.98 LPIPS $\downarrow$ E2VID [22, 36]+NeRF [8] 0.13 0.24 0.16 0.11 0.28 0.16 0.15 0.18 EventNeRF [12] 0.06 0.07 0.05 0.05 0.13 0.07 0.03 0.07 E2VID [22, 36]+3DGS [9] 0.12 0.17 0.16 0.11 0.19 0.16 0.12 0.15 Ours (Exposure Events) 0.01 0.02 0.01 0.01 0.05 0.01 0.04 0.02
Table 3: Quantitative Comparisons of Neural Reconstruction in Real World.
@@ 4.4 Quantitative Comparison
Figure 3: Qualitative comparison of appearance reconstruction on the real-world EME-3D dataset. The E2VID + NeRF method successfully reconstructs the overall scene but lacks fine local details, such as the tortoise’s shell texture. EventNeRF exhibits noticeable artifacts in both the car and the tortoise. In contrast, our proposed E-3DGS, in both fast and high-quality reconstruction modes, preserves sharper and more consistent structures while maintaining cleaner backgrounds.
@@
Figure 4: Qualitative comparison of geometry reconstruction on the real-world EME-3D dataset. Both E2VID+NeRF and EventNeRF, which are based on NeRF, struggle to separate the foreground from the background and are affected by noticeable noise. In contrast, the physics-based 3DGS method handles geometry reconstruction more effectively. Compared to E2VID+3DGS, our E-3DGS excels in capturing high-frequency spatial details, such as the gap between the crab’s body and the base.
Figure 5: Comparison of standard frame-based cameras and grayscale images generated from exposure events under extreme lighting conditions. The first column shows grayscale images derived from EventNeRF [12] ground truth RGB images with foreground masks extracted using EfficientSAM [33], serving as references under favorable lighting conditions. The second column presents simulated frame-based images under low light and overexposure. The third column presents grayscale images mapped from exposure events, derived from the frame-based images using the method described in Section 4.1. These images exhibit improved performance in low-light conditions and demonstrate the capability to recover details lost in overexposed regions, leveraging the high dynamic range and temporal resolution of event cameras.
4.5 Qualitative Comparison
@@
We qualitatively compare the appearance and geometric reconstruction of five sequences from the real-world EME-3D dataset by visualizing the rendered results and depth maps. Fig. 3 shows that our E-3DGS, in both fast and high-quality reconstruction modes, preserves sharper, more consistent structures and cleaner backgrounds compared to EventNeRF [12]. When using motion events, E-3DGS avoids artifacts seen in EventNeRF’s rendering, such as with the car ($1^{st}$ row) and the tortoise ($4^{th}$ row). Our results also exhibit higher contrast and sharper spatial details, particularly in regions with highlights and reflections, like the camera ($3^{rd}$ row) and the tortoise ($4^{th}$ row). Compared to E2VID [22] + NeRF, E-3DGS produces clearer local details, such as the tortoise’s shell texture, which is completely lost in the $4^{th}$ row of E2VID + NeRF. Furthermore, compared to E2VID + 3DGS, E-3DGS integrates exposure event information and reconstructs the overall grayscale tone and brightness more accurately. For instance, in the $3^{rd}$ row, E2VID + 3DGS shows white artifacts on the black camera body, while E-3DGS renders a clean, pure black camera body.
As shown in Fig. 4, E-3DGS achieves cleaner backgrounds, fewer artifacts, and richer spatial details in geometric reconstruction compared to both EventNeRF and NeRF integrated with E2VID. EventNeRF’s depth maps struggle to distinguish foreground from background and suffer from noticeable noise, such as with the car ($1^{st}$ row), motorcycle ($2^{nd}$ row), tortoise ($4^{th}$ row), and crab ($5^{th}$ row). While the depth map quality of E2VID + 3DGS is comparable to E-3DGS, E-3DGS excels in high-frequency spatial details. For example, in the crab sample, E-3DGS successfully reconstructs the gap between the crab’s body and the base, while E2VID + 3DGS confuses the subject with the background. Furthermore, the depth maps obtained in the fast reconstruction mode exhibit contours that align more closely with the actual geometry of the objects, providing a significant advantage when converting the 3DGS representation into mesh structures.
Previous qualitative and quantitative experiments have demonstrated the photometric accuracy advantage of our High-Quality Reconstruction Mode in utilizing exposure events for 3D reconstruction compared to motion events. While capturing exposure events may seem to compromise the event camera’s ability to handle fast-moving scenes, it retains its inherent advantage over standard frame-based cameras under extreme lighting conditions. As shown in Fig. 5, we simulate grayscale images mapped from exposure events under low-light and overexposure conditions using the method described in Section 4.1. Under low-light conditions, standard frame-based cameras lose the ability to effectively capture contour and fine structural details due to their limited dynamic range and reduced signal-to-noise ratio. In contrast, grayscale images mapped from exposure events exhibit a significantly higher dynamic range, enabling better detail preservation. Furthermore, in cases of overexposure, RGB cameras struggle with pixel saturation, where the sensor reaches its maximum capacity to record light intensity. This saturation leads to a loss of crucial details in high-brightness regions, such as reflective surfaces or intense light sources, making these areas appear as uniform white with no discernible features. However, since event cameras record brightness changes rather than absolute intensity, the grayscale images mapped from exposure events can effectively retain details in overexposed areas, showcasing their robustness in extreme illumination scenarios.
4.6 Ablations
@@
Figure 6: Qualitative comparison of different initialization methods. The results demonstrate the impact of initialization on the performance of motion event-based and exposure-enhanced 3DGS approaches. When using random initialization, the motion event-based 3DGS method produces noticeable artifacts and lacks fine details, as observed in the top-left image compared to the bottom-left. In contrast, the E-3DGS method, which integrates spatially rich exposure events, shows significantly improved robustness to initialization. As seen in the top-right and bottom-right images, the rendering results remain consistent and free of significant artifacts, regardless of the initialization method.
Table 4: Impact of Initialization Methods on Fast and HQ Modes.
@@
Table 5: Ablation Study on Different Initialization Methods (Fast Mode).
@@
Point Cloud Initialization. In Tab. 4, we explore the impact of different point cloud initialization methods on rendering performance in event-based 3DGS. Compared to random initialization, using a sparse point cloud from SfM [13] significantly improves the rendering accuracy when using motion events, with PSNR elevating to $25.22dB$ from $15.72dB$. However, in high-quality reconstruction mode, the PSNR is not sensitive to whether the initialization is from SfM or random point clouds ($33.82dB$ vs. $33.92dB$). As shown in Fig. 6, we visualize the effects of different initialization methods on event-based 3DGS rendering. When using random initialization, motion event-based 3DGS produces noticeable artifacts and limited details (top-left vs. bottom-left). In contrast, after integrating the spatially rich exposure events, E-3DGS becomes more robust to the initialization method, with no significant differences in rendering results (top-right vs. bottom-right). To further illustrate the advantages of incorporating exposure events in 3D reconstruction, we compare NeRF-based and event-to-video-based point cloud initialization methods in Tab. 5. Initializing the point cloud using EventNeRF-rendered results, compared to random initialization, provides a modest improvement in PSNR ($17.13dB$ vs. $15.72dB$), though it requires additional training for EventNeRF with limited gains. Using E2VID [22] to convert events into images followed by SfM for point cloud initialization further enhances accuracy ($23.47dB$ vs. $15.72dB$), but this process also involves additional learning-based computation. By fully leveraging the rich spatial cues embedded in exposure events, the proposed exposure event temporal-spatial mapping combined with SfM initialization yields the greatest improvement in reconstruction quality ($25.22dB$ vs. $15.72dB$). This certifies the critical role of exposure events in enhancing the quality of explicit 3D reconstruction from events.
Figure 7: Effects of integrating motion events with exposure events on 3D reconstruction for the real-world car sequence. Incorporating exposure event frames provides more accurate spatial information to constrain the 3D Gaussian splatting optimization, leading to steady improvements in PSNR and SSIM. Even a small number of exposure frames enhances reconstruction quality compared to using only motion events, highlighting the balanced hybrid mode’s effectiveness in fast-motion scenarios.
Table 6: Quantitative Results of Novel View Reconstruction.
@@
Novel View Synthesis. In this task, we train the model using the odd-numbered frames (Given) and test its ability to generate novel views using the even-numbered frames (Novel). As shown in Tab. 6, E-3DGS has excellent novel view synthesis performance in both fast reconstruction mode and high-quality reconstruction mode. On average, E-3DGS using motion events exhibits only a $0.05dB$ drop in PSNR for novel view synthesis compared to the given frames ($23.04dB$ vs. $23.09dB$). Similarly, E-3DGS using exposure events shows a $0.06dB$ PSNR drop for novel views compared to the given frames ($24.83dB$ vs. $24.89dB$). This demonstrates E-3DGS’s strong capabilities in both 3D reconstruction and novel view synthesis tasks.
5 Conclusion and Discussions
@@
Conclusion. In this work, we have presented E-3DGS, the first method to integrate 3DGS with exposure events for explicit scene reconstruction using a single pure high-resolution event sensor. By partitioning events into motion and exposure categories, E-3DGS leverages motion events for efficient 3D Gaussian Splatting (3DGS) reconstruction and utilizes exposure events to enhance reconstruction quality through high-resolution temporal-to-intensity mapping. Our framework is versatile and capable of operating on motion/exposure events alone or adopting a hybrid mode that balances reconstruction quality and speed by combining initial exposure events with high-speed motion events. On the EventNeRF dataset, we validate the feasibility of integrating a transmittance adjustment device with event cameras for high-quality 3D reconstruction. Under extreme lighting conditions, such as low illumination and overexposure, conventional frame-based cameras struggle with detail loss due to limited dynamic range. In contrast, the hardware-level adjustment preserves the high dynamic range advantage of event cameras, making them more effective for imaging in challenging environments. Additionally, we introduced EME-3D, a real-world dataset with exposure and motion events, camera calibration parameters, and sparse point clouds to support further research in this domain. E-3DGS outperforms event-based NeRF in both reconstruction speed and quality while being more cost-effective than methods that require both event and RGB data. By fully utilizing both motion and exposure events, E-3DGS sets a new benchmark for event-based 3D reconstruction, demonstrating robust performance in challenging conditions with reduced hardware demands.
Limitations. Despite the promising results, our current work has several limitations. Firstly, our method currently reconstructs grayscale scenes, lacking color information inherent in traditional RGB-based approaches. This limits its applicability in domains where color is essential for interpretation or downstream tasks. Secondly, we have not evaluated performance under extremely high event rates, where event camera saturation could degrade results, particularly in scenarios like large scenes under bright sunlight. Our system currently lacks specific mechanisms for handling such conditions. Thirdly, experiments were confined to smaller datasets, and the method’s scalability to large-scale environments remains untested.
Future work. Future work should prioritize addressing these limitations. Key directions include developing strategies for robustness against high event rates, such as adaptive event processing or integrating saturation models. Extending and evaluating E-3DGS on large-scale scenes is another critical goal, likely requiring algorithmic optimizations for efficiency and enhanced robustness to diverse real-world conditions. Crucially, we plan to explore the integration of color information, potentially by leveraging color event cameras and developing algorithms capable of reconstructing colored 3D scenes from their sparse, asynchronous output. This would significantly broaden the applicability of our event-based reconstruction framework. Furthermore, we intend to investigate advanced fusion algorithms and hardware optimizations to further exploit the complementary strengths of both exposure events and motion events, potentially leading to hybrid systems effective across an even wider range of conditions.
Funding. Natural Science Foundation of Zhejiang Province (Grant No. LZ24F050003); National Natural Science Foundation of China (Grant Nos. 12174341 and 62473139).
Disclosures. The authors declare that there are no conflicts of interest related to this article.
Data availability. The data underlying the results presented in this paper will be publicly available on GitHub at https://github.com/MasterHow/E-3DGS.